OpenStudy (anonymous):

A ball is dropped from a height of 41m. Each time it bounces it reaches 85% of it's previous height. When it hits the ground for the 26th time it has traveled Answerm. (Give answer to 2 decimal places)

OpenStudy (johnweldon1993):

Hmm lets see, does this want total distance traveled from the beginning? Or just how high the ball is in difference to the beginning. I believe the prior, So the common ratio is going to be .85 a1 (the beginning height) will be 41 m So the sum of the total distance traveled will be solved using \[\large S_{26} = \frac{a_1(1 - r^n)}{1 - r}\] so plugging in everything you have \[\large S_{26} = \frac{41(1 - (.85)^{26})}{1 - .85}\] \[\large S_{26} = \frac{41(1 - (0.01461813837))}{.15}\] \[\large S_{26} = \frac{41(0.98538186163)}{.15}\] \[\large S_{26} = \frac{40.4006563268}{.15}\] \[\large \S_{26} = 269.34\] This will be the TOTAL distance that ball has travelled from the drop, to the 26th bounce

OpenStudy (kropot72):

The distance traveled between two successive bounces will be twice the height reached. Therefore the sequence in terms of distance traveled is of the form: \[a,\ 2ar,\ 2ar ^{2},\ 2ar ^{3},\ 2ar ^{4},.........,\ 2ar ^{25}\] The total distance traveled is given by: \[\large \frac{2\times41(1-0.85^{26})}{1-0.85}-\frac{41}{1}\] Note that 41 is subtracted because the distance of 41 m is traveled only once.

OpenStudy (anonymous):

thanks guys! :D

OpenStudy (kropot72):

So I calculate the total distance traveled to be 497.7 m.

OpenStudy (anonymous):

what why are the 2 answers different o.o

OpenStudy (kropot72):

Ask @johnweldon1993 to list the terms of his sequence in distances traveled (as I have done). Then compare the two sequences.